OpenStudy (anonymous):

How many different committes can be formed from 7 teachers and 41 students if the committee consists of 2 teachers and 2 students

OpenStudy (anonymous):

the short answer is 7 choose 2 times 41 choose 2

OpenStudy (anonymous):

sometimes written as \[\binom{7}{2}\times \binom{41}{2}\] or \[_7C2\times _{41}C_2\] or even \[^7C_2\times ^{41}C_2\]

OpenStudy (anonymous):

do you know how to compute these numbers?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

they are both computed similarly \[\binom{7}{2}=\frac{7\times 6}{2}\] and \[\binom{41}{2}=\frac{41\times 40}{2}\]

OpenStudy (anonymous):

oh so the answer is 17220?

OpenStudy (anonymous):

i don't know, i didn't do it

OpenStudy (anonymous):

yeah, looks good http://www.wolframalpha.com/input/?i=%287+choose+2%29*%2841+choose+2%29

OpenStudy (anonymous):

sweet thanks

OpenStudy (anonymous):

yw